the school that stefan goes to is selling tickets to a choral performance. on the first day of ticket sales the school sold 3 senior citizen tickets and 1 child ticket for a total of $38. the school took in $52 on the second day by seelling 3 senior citizen and 2 child tickets. find the price of a senior citizen ticket and the prics of a child ticket.
Answer by jorel1380(3719)

3x+y=38
3x+2y=52
------------
3x+2y-3x-y=52-38=14
y=14
x=8
Senior tickets are $8. Children's are $14..